Please can you help me find x
((( 8x³/2x – 3x – 2 = 0 )))

Answer by stanbon(75887)

8x³/2x – 3x – 2 = 0
If the 2x is the denominator of the 8x^3 you get:
4x^2-4x-2=0
Using the quadratic formula you get:
x=[4+-sqrt(16-4*4*(-2)]/8
x=[4+-sqrt(48)]/8
x=[4+-4sqrt3]/8
x=[1+-sqrt3]/2
x=(1/2)+(1/2)sqrt3 or x=(1/2)-(1/2)sqrt3
Cheers,
Stan H.